OpenStudy (anonymous):

i give medals! A pencil is rolled off a table of height 1.25 m. If it has a horizontal speed of 3.1 m/s, how long does it take the pencil to reach the ground? A. 0.51 s B. 1.12 s C. 0.40 s D. 2.48 s

OpenStudy (anonymous):

The vertical motion is independent of the horizontal motion, so the change in vertical distance is y = 1.25 = (1/2) g t^2.

OpenStudy (anonymous):

i dont think im working it out correctly heres my work; 1.25=.5(9.8)(t^2) 1.25=4.9(3.1^2) 1.25=47.1 =.02 ? @douglaswinslowcooper